Python Coding:

Write a computer program that stores and tracks information about high school students. The first piece of information is what grade the student is in. Since this is for high school, the available values are 9, 10, 11, and 12. We also want to track the student’s GPA. Examples of GPAs are 2.2, 2.6, and 4.0. Finally, we want to track the letter grade that the student got on his or her final exam. These values are A, B, C, D, and F.

Using the programming language(s) that you select, write a program that creates variables to store these values. Devise with good names for these variables. In addition to creating the variables, set the values as follows:

Grade 9
GPA 3.5
Final grade B

When finished, your program should run without errors. It won’t do anything, because we have not included anything to be printed on the screen. So, when you run the program it should just terminate without errors.
Answer #1

IF YOU HAVE ANY DOUBTS COMMENT BELOW I WILL BE THERE TO HELP YOU

ANSWER:

CODE:

id_nums = []
grades = []
gpas = []
final_grades = []

def student_data(id_num,grade,gpa,final_grade):
id_nums.append(id_num)
grades.append(grade)
gpas.append(gpa)
final_grades.append(final_grade)
  
  
while(1):
print "Do you want enter the student data or view?"
choice = raw_input("Enter you choice (E - Enter/V - View): ")
if(choice == 'E'):
print "Valid data specifications: "
print "1). Id must be of 6 digits"
print "2). Grade must be in between 9-12 both included"
print "3). GPA must be less than 10.00"
print "4). Final grade must be one of A, B, C, D or F"
  
id_num = int(raw_input("Enter student Id number: "))
grade = int(raw_input("Enter grade: "))
gpa = float(raw_input("Enter gpa: "))
final_grade = raw_input("Enter final grade: ")
  
if((id_num<=99999 or id_num>=1000000) or (grade<9 and grade>12) or (gpa<0.0 and gpa>10.00) or (final_grade not in ('A','B','C','D','F'))):
print "Enter valid data as specified "
continue
else:
student_data(id_num,grade,gpa,final_grade)
  
elif(choice == 'V'):
print "Enter six digit Id number only.."
id_num = int(raw_input("Enter student Id number to serach: "))
if(id_num>99999 and id_num<1000000):
index = id_nums.index(id_num)
print "Student grade: ",grades[index]
print "Student GPA: ",gpas[index]
print "Student final grade: ",final_grades[index]
else:
print "Enter valid data.. "
continue

OR

CODE:

source code:

RATE THUMBSUP PLEASE